# Direct Density Ratio Optimization: A Statistically Consistent Approach to Aligning Large Language Models

### v1

@@ 1 Introduction
However, these existing methods rely on the assumption that human preferences follow a specific model (e.g., the Bradley-Terry model (Bradley & Terry, 1952)). This assumption may not accurately capture the complexity of real human preferences (Wu et al., 2022), leading to a fundamental issue of a lack of statistical consistency. Statistical consistency refers to the property that, as the amount of data increases, the learned model converges to the true optimal model (Vapnik, 1998; Mohri, 2018). Without statistical consistency, increasing the amount of data does not guarantee performance improvement, raising concerns about the model’s reliability and safety. More concretely, even if one perfectly optimizes the loss function based on the Bradley-Terry model, it does not necessarily guarantee the accurate learning of the true underlying human preferences.
@@ 3.2 DDRO: The Proposed Method
The Bregman divergence framework that we employed for DDRO was originally established in (Sugiyama et al., 2012b) and is one of the most general framework for density ratio estimation. This framework encompasses various density ratio estimation methods depending on the choice of Bregman divergence (i.e., the convex function $f$) and how the density ratio model is parameterized. Notable examples include Least Squares Importance Fitting (LSIF) (Kanamori et al., 2009), Kernel Mean Matching (KMM) (Gretton et al., 2008), and Kullback-Leibler Importance Estimation Procedure (KLIEP) (Sugiyama et al., 2007b). Basically, we may employ any density ratio estimator for the preference optimization. However, we employ the logistic loss for our implementation due to its simplicity.
@@ 4.1 Estimation error bound
where $(\sigma_{i})_{i=1}^{n}$ is an i.i.d. Rademacher sequence, that is, $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$ (see Mohri et al. (2012) for a reference). It is known that the Rademacher complexity measures the variance of the empirical risk minimizer that reflects the complexity of the model. Let $\tilde{p}=\frac{1}{1-t}\frac{p_{\text{ref}}}{p}-\frac{t}{1-t}$ for $p\in\mathcal{H}$, and let $h^{\pm}:\mathbb{R}\to\mathbb{R}$ be $h^{+}(u)=-f\left(u\right)+uf^{\prime}\left(u\right)$ and $h^{-}(u)=-f^{\prime}\left(u\right)$. Then, we have the following theorem.
@@
Let $\hat{\theta}=\arg\min_{\theta\in\Theta}\mathcal{L}_{\mathrm{DDRO}}(\theta)$ be the optimal solution of the DDRO loss function with $\gamma=0$. We assume that $\|\tilde{p}/p\|_{\infty}\in I$ for all $p\in\mathcal{H}$ for some open bounded interval $I\subset\mathbb{R}$. We also assume that $h^{+}$ and $h^{-}$ are Lipschitz continuous on the interval $I$ with a Lipschitz constant $\mathrm{Lip}(h)$, $m_{+}:=\min_{x\in\mathcal{X},y\in\mathrm{supp}(p^{+}(\cdot\mid x))}p^{+}(y\mid x)>0$, and $\mu:=\inf_{s\in I}f^{\prime\prime}(s)>0$. Then, the estimation error of $p_{\hat{\theta}}$ with respect to the true preferred distribution $p^{+}$ satisfies:
@@
where the expectation is taken over the realization of the supervised data $\mathcal{D}_{\pm}$.
The proof is given in Appendix C. Since usual Rademacher complexity bounds yield $\mathcal{R}_{n}(\mathcal{H})=\mathcal{O}(1/\sqrt{n})$ (Mohri, 2018), the right hand side can be evaluated as $\mathcal{O}\left(\frac{1}{\sqrt{\min\{n^{+},n^{-}\}}}+\inf_{\theta\in\Theta}\mathcal{L}_{\mathrm{Breg}}(g_{\theta})\right)$. The term $\inf_{\theta\in\Theta}\mathcal{L}_{\mathrm{Breg}}(g_{\theta})$ represents how far the model is misspecified from the true distribution. If the true preferred data distribution is included in the model, this term vanishes. Hence, we see that the bound represents the bias and variance trade-off to estimate an appropriately aligned model. As the data size increases, DDRO becomes more accurate even though we are not putting any assumption on the human preference model.
@@ Acknowledgements
Theorem C.1 (Theorem 4.1 restated).
Let $\hat{\theta}=\arg\min_{\theta\in\Theta}\mathcal{L}_{\mathrm{DDRO}}(\theta)$ be the optimal solution of the DDRO loss function with $\gamma=0$. Let $I$ be an open bounded interval and we assume that $\|\tilde{p}/p\|_{\infty}\in I$ for all $p\in\mathcal{H}$. We also assume that $h^{+}$ and $h^{-}$ is Lipschitz continuous on the interval $I$ with a Lipschitz constant $\mathrm{Lip}(h)$, $m_{+}:=\min_{x\in\mathcal{X},y\in\mathrm{supp}(p^{+}(\cdot\mid x))}p^{+}(y\mid x)>0$, and $\mu:=\inf_{s\in I}f^{\prime\prime}(s)>0$. Then, the estimation error of $p_{\hat{\theta}}$ with respect to the true preferred distribution $p^{+}$ satisfies:
$\mathbb{E}_{\mathcal{D_{\pm}}}\left[\left\|p_{\hat{\theta}}-p^{+}\right\|_{L^{2}(p^{+})}^{2}\right]\leq\frac{2(1-t)^{2}}{t^{2}m_{+}^{2}\mu}\left[4\mathrm{Lip}(h)(\mathcal{R}_{n^{+}}(\mathcal{H})+\mathcal{R}_{n^{-}}(\mathcal{H}))+\inf_{\theta\in\Theta}\mathcal{L}_{\mathrm{Breg}}(g_{\theta})\right],$ (37)
where the expectation is taken over the realization of the supervised data $\mathcal{D}^{\pm}$.

### v2

@@ 1 Introduction
However, these existing methods rely on the assumption that human preferences follow a specific model (e.g., the Bradley-Terry model (Bradley & Terry, 1952)). This assumption may not accurately capture the complexity of real human preferences (Wu et al., 2022), leading to a fundamental issue of a lack of statistical consistency. Statistical consistency refers to the property that, as the amount of data increases, the learned model converges to the true optimal model (Vapnik, 1998; Mohri et al., 2012). Without statistical consistency, increasing the amount of data does not guarantee performance improvement, raising concerns about the model’s reliability and safety. More concretely, even if one perfectly optimizes the loss function based on the Bradley-Terry model, it does not necessarily guarantee the accurate learning of the true underlying human preferences.
@@ 3.2 DDRO: The Proposed Method
The Bregman divergence framework that we employed for DDRO was originally established in (Sugiyama et al., 2012b) and is one of the most general framework for density ratio estimation. This framework encompasses various density ratio estimation methods depending on the choice of Bregman divergence (i.e., the convex function $f$) and how the density ratio model is parameterized. Notable examples include Least Squares Importance Fitting (LSIF) (Kanamori et al., 2009), Kernel Mean Matching (KMM) (Gretton et al., 2009), and Kullback-Leibler Importance Estimation Procedure (KLIEP) (Sugiyama et al., 2007b). Basically, we may employ any density ratio estimator for the preference optimization. However, we employ the logistic loss for our implementation due to its simplicity.
@@ 4.1 Estimation error bound
where $(\sigma_{i})_{i=1}^{n}$ is an i.i.d. Rademacher sequence, that is, $P(\sigma_{i}=1)=P(\sigma_{i}=-1)=1/2$ (see Mohri et al. (2012) for a reference). It is known that the Rademacher complexity measures the variance of the empirical risk minimizer that reflects the complexity of the model. Let $\tilde{p}=\frac{1}{1-t}p_{\text{ref}}-\frac{t}{1-t}p$ for $p\in\mathcal{H}$, and let $h^{\pm}:\mathbb{R}\to\mathbb{R}$ be $h^{+}(p)=-f\left(\tilde{p}/p\right)+\left(\tilde{p}/p\right)f^{\prime}\left(\tilde{p}/p\right)$ and $h^{-}(p)=-f^{\prime}\left(\tilde{p}/p\right)$. Then, we have the following theorem.
@@
Let $\hat{\theta}=\arg\min_{\theta\in\Theta}\mathcal{L}_{\mathrm{DDRO}}(\theta)$ be the optimal solution of the DDRO loss function with $\gamma=0$. We assume that there exists some open bounded interval $I\subset\mathbb{R}$ such that $\tilde{p}(y|x)/p(y|x)\in I$ for all $p\in\mathcal{H}$, $x\in\mathcal{X}$, and $y\in\mathcal{Y}$. We also assume that $h^{+}$ and $h^{-}$ are Lipschitz continuous on the interval $I$ with a Lipschitz constant $\mathrm{Lip}(h)$. Then, the estimation error of $p_{\hat{\theta}}$ with respect to the true preferred distribution $p^{+}$ satisfies:
@@
where the expectation is taken over the realization of the supervised data $\mathcal{D}_{\pm}$, $m_{+}:=\min_{x\in\mathcal{X},y\in\mathrm{supp}(p^{+}(\cdot\mid x))}p^{+}(y\mid x)>0$, and $\mu:=\inf_{s\in I}f^{\prime\prime}(s)>0$.
The proof is given in Appendix C. Since usual Rademacher complexity bounds yield $\mathcal{R}_{n}(\mathcal{H})=\mathcal{O}(1/\sqrt{n})$ (Mohri et al., 2012), the right hand side can be evaluated as $\mathcal{O}\left(\frac{1}{\sqrt{\min\{n^{+},n^{-}\}}}+\inf_{\theta\in\Theta}\mathcal{L}_{\mathrm{Breg}}(g_{\theta})\right)$. The term $\inf_{\theta\in\Theta}\mathcal{L}_{\mathrm{Breg}}(g_{\theta})$ represents how far the model is misspecified from the true distribution. If the true preferred data distribution is included in the model, this term vanishes. Hence, we see that the bound represents the bias and variance trade-off to estimate an appropriately aligned model. As the data size increases, DDRO becomes more accurate even though we are not putting any assumption on the human preference model.